Operation Summary for June 28, 2010 - June 30, 2010
Monday June 28
The day shift began with Operations monitoring the Tevatron (TeV) store 7931 (93.47E30) and the Recycler stash (218.4E10), with the Antiproton Source (Pbar) stacking (27.3E10), with the Linac, Booster and the Main Injector (MI) operating normally, with the NuMI experiment off due to an access, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the evening shift, the TeV quenched at 4:56 PM during the squeeze. Recovery and TeV studies followed. At 6:57 PM, the stash of 75E10 antiprotons was lost due to a bulk power supply (R:C621) trip.
Tuesday June 29
The midnight shift began with a TeV person conducting studies, with Operations monitoring the stash (108.9E10), with Pbar stacking (8.5E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations put the TeV into shot setup at 3:43 AM and established store 7938 at 4:49 AM with an average initial luminosity (AIL) of 184.4E30.
During the day shift, Operations reported at 10:33 AM that a MI quadrupole power supply (I:Q847) tripped off due to a bad controller. Operations had the MI beam restored by 3:08 PM.
During the evening shift, Operations terminated store 7938 at 11:56 PM.
Wednesday June 30
The midnight shift began with Operations monitoring the stash (356.3E10), with Pbar stacking (25.3E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations put the TeV into shot setup and established store 7939 at 1:16 AM with an AIL of 345E30.
The Plans for Wednesday
Comments and Suggestions
|last modified 06/30/2010 email Fermilab|
|Security, Privacy, Legal|